2. A solution was produced by dissolving 3.75 g of a nonvolatile solute in 95g of


acetone. The boiling point of pure acetone was observed to be 55.95 ºC, while the boiling


point of the solution was 56.50ºC. If the boiling point for acetone is 1.71ºC/m, what is the


approximate molar mass of solute?


Step 1: First compute the molality of the boiling point equation.


Step 2. Then, from the definition of molality, compute the number of moles


solute, n(solute), in the sample.


Step 3. Solve for the molar mass.



Please solve it like this




STEP1: SOLUTION



STEP2: SOLUTION



STEP3: SOLUTION




THANK YOUUU

1
Expert's answer
2022-03-27T10:30:44-0400

deltaT = Cm x Kb

deltaT = 56.5 - 55.95 = 0.55 degrees

Cm = deltaT / Kb = 0.55 / 1.71 = 0.32 m

The formula for molality is m = moles of solute / kilograms of solvent.

Moles of solute = Сm x kilograms of solvent = 0.32 x 0.095 = 0.03 mol

n = m / M

Msolute = m / n = 3.75 / 0.03 = 125 g/mol